Can you please help me solve using the elimination method?
In a recent game, Charlie scored 42 points on a combination of 18 two and three point baskets. How many shots of each type were made?

Answer by checkley71(8403)

2x+3(18-x)=42
2x+54-3x=42
-x=-12
x=-12/-1
x=12 solution
proof
2*12+3(18-12)=42
24+54-36=42
78-36=42
42=42